# Response to “Response to ‘Comment on Theoretical examination of QED Hamiltonian in relativistic molecular orbital theory”’ [J. Chem. Phys. 160, 187102 (2024)] Wenjian Liu

### v5

@@
Inoue and coworkers’ Response[1] to my Comment[2] on their Paper I[3] clarifies some points,
@@
As far as the nonrelativistic limit of is concerned, Inoue and coworkers
@@
As emphasized from its introduction[4] (see also Ref. 6 for a thorough discussion),
@@
It is textbook knowledge that both the VP and ESE have to be regularized/renormalized before they can actually be
@@
Speaking of “QED Hamiltonian” in the absence of VP-ESE, the genuine QED effect[4], is hardly meaningful.
@@
To resolve this apparent self-contradiction, they separated strictly the use of the CNC/CCC from that of the
@@
In other words, the contraction over the vacuum
@@
Since the Hamiltonian misses by construction the genuine QED effect, it had better go back to its original name,

### v6

@@
The Response of Inoue and coworkers[1] to my Comment[2] on their Paper I[3] clarifies some points,
@@
As far as the nonrelativistic limit of (4) is concerned, Inoue and coworkers
@@
As emphasized from its advent[4] (see also Ref. 6 for a thorough discussion),
@@
It is then textbook knowledge that both the VP and ESE have to be regularized/renormalized before they can actually be
@@
Speaking of “QED Hamiltonian” in the absence of the VP-ESE (genuine QED effect[4]) is hardly meaningful.
@@
To resolve this apparent self-contradiction, they separated the CNC/CCC from the
@@
In other words, being imposed from the outset, the CCC over the vacuum
@@
It can also be shown that the CCC arises naturally from the averaging of
the Hamiltonians for electrons and positrons moving in the same external field (which is a must, as shown in Appendix A).
Since the Hamiltonian misses by construction the genuine QED effect[4], it had better go back to its original name,
@@ Acknowledgement
Appendix A Hamiltonian quantized with charge conjugated fields
The charge conjugation is defined as
which transforms the one-electron Dirac equation
to that of a positron
Likewise, the four-component Dirac field operator (in the particle-hole picture)
will be transformed by charge conjugation to
where the first and second terms annihilate a positive-energy positron (NB: ) and create a positive-energy electron, respectively,
in accordance with the fact that the first and second terms of (31)
annihilate a positive-energy electron and create a positive-energy positron, respectively. That is,
(33) would be the starting quantized Dirac field, had we lived in the world of antiparticles.
Both (31) and
(33) are associated with the vacuum .
To expedite algebraic manipulations, we can rewrite them as[6]
by taking (Dirac sea of electrons)
and (Dirac sea of positrons)
as the vacua, respectively.
With the above background, we can calculate
which is to be compared to the usual second-quantized Dirac operator
Since charge conjugation is an inherent symmetry, (39)/(40) and
(36)/(37) should be averaged with an equal weight, leading to
where is just the CCC (8) and arises here naturally from the averaging process.
The two-body operator can be calculated in the same way, viz.,
which should be compared to the usual two-body operator
The average of and leads to
It can be seen that the sum of (38) and (50) is just (4) along with the CCC. In particular,
in sharp contrast to the -potential in Eq. (50),
the second term of (46) or that of (48) is divergent and cannot be rgularized/renormalized.
As such, the averaging of the Hamiltonian for electrons and the Hamiltonian for positrons (moving in the
same external field as electrons)
is a must rather than merely a formal step. Note in passing that
the constant (51) is different from the second term of Eq. (7) (which is more symmetric due to
the direct use of the CCC). However, the difference does doe not matter at all, for such constants will be renormalized away.
The above manipulation reveals that the genuine QED Hamiltonian (4), especially the Q-potential in Eq. (50)
[or Eq. (6)/(13)],
does stem from the symmetric treatment of electrons and positrons, as dictated by charge conjugation symmetry of relativistic fermions.
In particular, Eqs. (37), (40), (46), and (48) all follow strictly the ACR (9).